A staff member mixed a 35% solution with a 10% solution. How many liters of each will she need to mix in order to obtain 450 liters of a 20% solution?
Answer by stanbon(75887)

A staff member mixed a 35% solution with a 10% solution. How many liters of each will she need to mix in order to obtain 450 liters of a 20% solution?
--------------------------
Equation:
0.35x + 0.10(450-x) = 0.20*450
--------------------
Multiply thru by 100 to get:
35x + 10(450-x) = 20*450
35x + 10*450 - 10x = 20*450
25x = 10*450
x = 10*18
x = 180 liters (amt. of 35% solution needed in the mixture)
---
450-180 = 270 liters (amt. of 10% solution needed in the mixture)
====================================================================
Cheers,
Stan H.